Two capacitors c1 and c2 are connected in parallel. The total charge of the capacitors is Q. What is the voltage on the system? What are the charges of the capacitors ?
Answer #1

Since C1 and C2 are in parallel , their parallel combination is given as

C = C1 + C2

Q = total charge

Voltage is given as

V = Q/C

V = Q/ (C1 + C2)

Charge on capacitor C1 is given as

Q1 = C1 V = C1 Q/ (C1 + C2)

Charge on capacitor C2 is given as

Q2 = C2 V = C2 Q/ (C1 + C2)